OpenStudy (angela210793):

suppose there are 100 guys... and each of them pick any number between 0 and 100 once everybody has picked a number....you add all of the numbers and multiply that by 2/3 so...if everybody picked 100....the average would be 100 and after multiplying by 2/3, you get 66.67... now, whoever's picked number is closest to the average multiplied by 2/3, wins the game and gets $100..... what no I should I choose to maximize my chance of winning?

OpenStudy (anonymous):

Just a stab in the dark, but if you multiplied each number by 2/3, and found the average of those. Wouldn't that be the best number to guess?

OpenStudy (angela210793):

Idk....I've no idea at all :/

OpenStudy (anonymous):

Use Bayes' Theorem!

OpenStudy (angela210793):

Bayes????????

OpenStudy (anonymous):

what is the most likely number that 100 guys would pick?

OpenStudy (anonymous):

1 ?

OpenStudy (angela210793):

Not getting it..........

OpenStudy (anonymous):

good! so if 100 guys pick one, then the result would be 2/3!

OpenStudy (anonymous):

so if you picked 1, then you would be winning

OpenStudy (angela210793):

come on....i don't think it is solved like this guy1=59 guy2=31 . . . . . . . . . . .guy100=3 Wht now?

OpenStudy (anonymous):

alright, try having all guys pick different numbers: guy 1 pick 1, guy 2 pick 2... guy 100 pick 100

OpenStudy (anonymous):

find\[\sum_{i=1}^{100}i\]

OpenStudy (anonymous):

this is \[\frac{100*(100+1)}{2}=10100/2 = 5050\]

OpenStudy (angela210793):

i did smth like